OpenStudy (anonymous):

What is the equation of the line that contains the points (3, –6) and (–3, 0)? A. y = x – 3 B. y = –x – 3 C. y = –x + 3 D. y = x + 3

OpenStudy (anonymous):

its D

OpenStudy (luigi0210):

Find the slope first: \[\LARGE \frac{y_{2}-y_{1}}{x_{2}-x_{1}}\] Then use: \[\LARGE y-y_{1}=m(x-x_{1})\] @sabrinamedina77 Don't give out direct answers

OpenStudy (anonymous):

Ok ,im new to this open study thing ,i just need tofinish a few assignments to catch up ,and thanks :)

OpenStudy (luigi0210):

Well \[\Huge \mathbb {\color{green}{WELCOME}}\] And besides, D was the wrong answer ._.